OpenStudy (anonymous):

Factor completely: 2x2 + 4x − 2 2(x2 + 2x − 1) 2(x2+ 2) 2x(x2 + 2x − 1) Prime

OpenStudy (anonymous):

2x^2+4x-2

OpenStudy (anonymous):

\[2x^{2}+4x-2=2.x^{2}+2.2x-2.1\] Nw take 2 common from all the expressions @harliedanielfrye

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

@harliedanielfrye split the expressions in such a way that they have a common factor you can write \[2x^{2}=2*x^{2}\]\[4x=2*2x\]\[2=2*1\]

OpenStudy (anonymous):

so u can write 2x^2 + 4x − 2=2*x^2+2.2x-2.1 @harliedanielfrye

OpenStudy (anonymous):

have u undrstd till here ??? @harliedanielfrye

OpenStudy (anonymous):

nope sorry i am really stupid

OpenStudy (anonymous):

suppose u need to factorise x*a+x*b-x*c u can easily factorise it by writing x(a+b-c) Here x=2 a=x^2 b=2x c=1 so 2.x^2+2.2x-2.1 =2(x^2+2x-1) Have u got it??? @harliedanielfrye

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good :) wtz ur age??? @harliedanielfrye

OpenStudy (anonymous):

15

OpenStudy (anonymous):

ook :) @harliedanielfrye wch country???

OpenStudy (anonymous):

r u asking where am i from?

OpenStudy (anonymous):

yep @harliedanielfrye

OpenStudy (anonymous):

thx @Curious9876